### 01 Dec 2023 v1

# Rethinking Detection Based Table Structure Recognition for Visually Rich Documents

Abstract
Table Structure Recognition (TSR) is a widely discussed task aiming at transforming unstructured table images into structured formats, such as HTML sequences. One type of popular solution is using detection models to detect components of a table, such as columns and rows, then applying a rule-based post-processing method to convert detection results into HTML sequences. However, existing detection-based studies often have the following limitations. First, these studies usually pay more attention to improving the detection performance, which does not necessarily lead to better performance regarding cell-level metrics, such as TEDS, after the post-processing. Second, some solutions over-simplify the problem and can miss some critical information. Lastly, even though some studies defined the problem to detect six components, including table, row, column, spanning cell, projected row header, and column header, to provide as much information as other types of solutions, these studies ignore the fact this problem definition is a multi-label detection because row, projected row header and column header can share identical bounding boxes. Besides these limitations, there is often a performance gap between two-stage and transformer-based detection models regarding the structure-only TEDS, even though they have similar performance regarding the COCO metrics. Therefore, in this study, we revisit the limitations of existing detection-based solutions, compare two-stage and transformer-based detection models, and identify the key design aspects for the success of a two-stage detection model for the TSR task, including the multi-class problem definition, the aspect ratio for anchor box generation, and the feature generation of the backbone network. Without proposing sophisticated new components, we applied simple methods to improve these aspects of the Cascade R-CNN model, achieved state-of-the-art performance, and improved the baseline Cascade R-CNN model by 19.32%, 11.56% and 14.77% regarding the structure-only TEDS on SciTSR, FinTabNet, and PubTables1M datasets.
keywords:
Table Structure Recognition, Information Extraction, Document Processing, Object Detection, Visually Rich Document Understanding
[…]
[inst1]organization=School of Electrical Engineering and Computer Science,addressline=University of Ottawa, city=Ottawa, postcode=K1N 6N5, state=ON, country=Canada
[…]
[inst2]organization=Lytica Inc.,addressline=555 Legget Drive, city=Ottawa, postcode=K2K 2X3, state=ON, country=Canada
1 Introduction
Figure 1: Generic workflow of transforming unstructured tabular data from document images into a structured format. Figure (a) is a sample document containing a table and other text areas. (b) represents the table detection step, and the output of this step is the table area. Note that the samples in (a) and (b) are only parts of the original results due to the page size limitation. (c) represents the step of table structure recognition. Following this study’s definition, its outputs are the detected Tables, Columns, Rows, Spanning Cells, and Headers. (d) represents the post-processing step reconstructing the complex table structure based on the detected table components. (e) is the OCR tool to extract text based on the reconstructed structure to fully transform the table image into a structured format. Notably, step (d) can also generate the cell boxes as the inputs of step (e).
(a) Four types of defined table components in the study in fernandes2023tablestrrec . This definition can infer spanning cells but cannot provide header information. Besides, the projected row headers are treated as rows, which can lead to a wrong structure.
(b) Four types of defined table components in the study in xiao2022efficient . This definition cannot provide header information, and the projected row headers are treated as rows, which can lead to a wrong structure.
(c) Six types of defined table components in the study in smock2022pubtables . The defined projected row headers in this sample share identical bounding boxes with two corresponding rows.
Figure 2: Different problem formulations for the detection-base TSR.
Figure 3: Four types of defined table components. In this sample, the defined column header shares an identical bounding box with a defined row.
Datasheets are widely used in business scenarios and on the internet because of their user-friendliness to human readers. However, these documents are usually not structured, which means that it is hard to extract useful information from them and further analyze the semantic meaning automatically. Besides the unstructured format of the document, tables in these documents, which are widely used to summarize critical information, often have very complex structures and layouts, making it challenging to interpret and analyze these tables. Therefore, there have been many studies of Table Detection (TD), Table Structure Recognition (TSR), Text Detection, and Optical Character Recognition (OCR) as steps of interpreting and analyzing tables from visually rich documents. Figure 1 shows a typical workflow of extracting tables from documents and transforming the extracted tables into an HTML format, consisting of a TD, a TSR, and an OCR steps. Considering that this study formulates the TSR as an Object Detection problem, a post-processing step is required after the TSR detection model, as shown in Figure 1(d). Since there have been many studies discussing Table Detection xiao2023revisiting ; xiao2023table ; fernandes2021tabledet and OCR tools mmocr2021 ; xiao2023multi , achieving promising results, we focus on the TSR in this study, which aims at transforming the tables from visually rich documents into structured formats.
TSR studies can roughly be categorized into three groups based on their problem formulations: detection-based models, image-to-sequence models, and graph-based models. Image-to-sequence models usually follow the encoder-decoder architecture and directly generate structured outputs, such as HTML sequences. Some image-to-sequence models visapp23namly ; ye2021pingan also integrate the OCR task into the model to make the model end-to-end without using extra OCR tools to extract text contents from the images. However, since these models use auto-regressive decoders, they often suffer from error accumulation shen2023divide , and their OCR capacity usually cannot generalize well because of the limitation of training data. On the other hand, graph-based models usually use segmentation or detection methods to extract table cells and treat extract table cells as nodes of a graph and further build the relation among the graph nodes. This graph-based definition makes it easier to deal with the scenarios in which table images are collected from the wild, such as rotated, distorted tables. However, graph-based models introduce extra complexity because of building graph models compared with detection-based models. By contrast, detection-based models provide an alternative way that is more straightforward to detect the table components directly and post-process the detection results with a deterministic rule-based method for reconstructing the table structure. However, the detection-based method can fail to deal with rotated and distorted samples. To sum up, these different types of approaches have their benefits and must be selected based on the application scenarios. This study focuses on the tables from well-formatted documents and proposes a detection-based solution.
There have been some studies siddiqui2019deeptabstr ; hashmi2021guided ; fernandes2023tablestrrec ; xiao2022efficient ; smock2022pubtables using detection models together with a post-processing method to solve the TSR task. However, existing studies either over-simplify the problem or define a multi-label detection problem, which is challenging for two-stage object detectors. For example, studies siddiqui2019deeptabstr ; hashmi2021guided formulate the problem as only detecting columns and rows, making them hard to deal with spanning cells. TableStrRec fernandes2023tablestrrec further extend the formulation by defining regular column, regular row, irregular column, and irregular row so that the spanning cell can be inferred from these four types of components, as shown in Figure 1(a). study in xiao2022efficient simplifies the formulation of study in smock2022pubtables , directly detecting table, column, row, and spanning cell, as shown in Figure 1(b). These three formulations treat the projected row header as a regular row and do not define headers, resulting in over-simplified table structures. By contrast, study in smock2022pubtables defines six types of components, including table, column, row, spanning cell, column header, and project row header, as shown in Figure 1(c), which can provide as much structure information as other types of TSR models. However, study in smock2022pubtables does not consider that some column headers and projected row headers can share identical bounding boxes with corresponding rows. For example, as shown in Figure 1(c), the bounding boxes of the project row header can also be classified as rows. Similarly, as shown in Figure 3, a column header’s bounding box is also the row’s bounding box. Therefore, the problem definition of study in smock2022pubtables is a multi-label detection problem. This multi-label definition is challenging for typical two-stage models, such as Faster R-CNN ren2015faster and Cascade R-CNN Cai2018CascadeRD , because the classifier for the classification task is for single-label classification problems, and the feature fed into the classifier are extracted from the backbone network instead of learned queries in Sparse R-CNN sun2021sparse and DETR carion2020end . Therefore, this study follows the definition of study in smock2022pubtables , analyzes the impact of multi-class definition to the Cascade R-CNN and Sparse R-CNN, which are typical two-stage and transformer-based detectors, and further refactors this multi-label definition by extending the overlapped samples to pseudo-classes to improve the performance of our proposed Cascade R-CNN based solution. Besides, the defined table components can have extremely large aspect ratios, and the ground truth bounding boxes contain a large number of white areas that cannot generate discriminate features for the classification and regression task, as shown in Figures 1(c) and  3. Therefore, suitable aspect ratios for the anchor generation of two-stage detectors and discriminate, long-range features from the backbone network are two critical aspects for the success of a two-stage detector for the TSR. Therefore, this study tunes the aspect ratio in the Region Proposal Network (RPN), applies deformable convolution, and introduces a Spatial Attention Module that can be inserted into the backbone network to enrich long-range and context features. To sum up, the contributions of this study are four-fold:
  1. 1.
    This study revisits popular detection-based TSR definitions and formulates the TSR problem as detecting six components: table, column, row, spanning cell, projected row header, and column header, following the definition of the study in smock2022pubtables . This formulation is a multi-label detection problem, which is challenging for two-stage detectors. Therefore, this study extends the formulation to conventional single-label detection problems by generating pseudo-classes.
  2. 2.
    This study identifies and analyzes critical design aspects for the success of a TSR model, including the aspect ratios in the anchor box generation and the discriminate feature generation, then further proposes TSRDet, a Cascade R-CNN based solution, by adjusting the ratio aspects applying the deformable convolution and introducing a spatial attention module to build the long-range dependencies and context information in the backbone network.
  3. 3.
    We conduct comprehensive experiments to evaluate our proposed solution on various datasets, including SciTSR chi2019complicated , FinTabNet zheng2020global , PubTabNet zhong2020image and PubTables1M smock2022pubtables with both detection metrics and cell-level metrics. The experimental results show that our proposed TSRDet can outperform state-of-the-art models in terms of both cell-level evaluation metrics and detection metrics.
  4. 4.
    We compare the two-stage and transformer-based detectors for the TSR problem, using the TSRDet and Sparse R-CNN as two examples, and further analyze the impact of the identified critical design aspects on the transformer-based detectors.
The rest of this paper is organized as follows: Section 2 discusses related studies, including studies in Object Detection Models and Table Structure Recognition models. Section 3 describes our proposed solution. Section 4 shows the experimental results and discusses the design aspects of the proposed method. At last, we draw our conclusion and possible directions in Section 5.
2 Related Work
2.1 Object Detection Models
Object Detection is a fundamental task that has been widely discussed in many studies. Since deep models have become the dominant solutions, we only discuss popular deep learning based models in this section. Based on different design perspectives, popular detection models can be categorized in different ways. One popular categorization of detectors based on the number of regression steps is to classify them into one- and two-stage detectors. Two-stage models, such as Faster R-CNN ren2015faster and Cascade R-CNN Cai2018CascadeRD , usually use a Region Proposal Network (RPN) to generate region proposals first and then feed the region proposals to the well-designed model to classify and regress the proposals. In the RPN network, one key parameter is the aspect ratio, which defines the height/width ratio when generating anchor boxes. Suitable aspect ratios are often close to the target objects’ height/width ratio, making the regression task easier and improving the model performance. In contrast, popular one-stage models, such as FCOS tian2019fcos , YOLO redmon2016you , and YOLO’s variants redmon2017yolo9000 ; redmon2018yolov3 ; bochkovskiy2020yolov4 , integrate the region proposal generation and other regression and classification components into a single network. For example, YOLO divides the images into grids first, then classifies the class of grid cells and directly predicts the bounding boxes and their confidences. The simple design of one-stage detectors leads to faster training and inference time compared with two-stage detectors.
On the other hand, some studies sun2021sparse ; zhang2023dense categorize the popular detectors from the perspective of Non-maximum Suppression (NMS), which is widely used to reduce the redundant predictions from the detectors. From this perspective, popular detectors can be categorized into end-to-end and none end-to-end models based on whether NMS is needed. DETR carion2020end is a typical end-to-end detector introducing transformer architecture vaswani2017attention , set prediction loss, and one-to-one label assignment to the object detection problem. Sparse R-CNN sun2021sparse further refactors the DETR model and proposes to use sparse learnable regional proposals to replace dense regional proposals and utilize a dynamic instance interactive head to regress and classify the proposals in an iterative manner. study in sun2021makes analyses the success of end-to-end detectors and argues that the one-to-one label assignment method in end-to-end detectors contributes to the success of the end-to-end models but is not sufficient to fully remove the NMS from the pipeline. This study further points out that the classification cost in the matching cost when applying one-to-one label assignment plays a key role in the success of these end-to-end models. study in zhang2023dense further analyzes combinations of the label assignment methods and queries and argues that sparse queries with one-to-one label assignment can degrade the recall, and dense queries with one-to-one label assignment are hard to optimize. To address these issues, study in zhang2023dense proposes a dense distinct queries (DDQ) method to select distinct queries from dense queries using a class-agnostic NMS, achieving promising precision and recall. It is worth mentioning that these end-to-end detector can easily be extended to none end-to-end solutions by adapting many-to-one label assignments hong2022dynamic and NMS.
2.2 Table Structure Recognition
There have been many studies Chi2019ComplicatedTS ; Adiga2019TableSR ; zheng2021global ; xiao2022efficient ; Schreiber2017DeepDeSRTDL ; liu2022neural discussing the TSR problem in recent years. As mentioned earlier in Section 1, we can roughly categorize these solutions into image-to-sequence, detection-based, and graph-based models. Image-to-sequence based models usually define the ground truth as structured sequences, such as HTML sequences, built on the transformer architecture vaswani2017attention , and follow an encoder-decoder architecture. For instance, TableMaster ye2021pingan is a typical image-to-sequence based model that can generate HTML sequences. More specifically, TableMaster follows the architecture of MASTER lu2021master , which is originally designed for the scene text generation following the transformer architecture vaswani2017attention , and further improved the encoder part by introducing a Multi-Aspect Global Context Attention. In contrast, TableMaster has two branches designed for the HTML sequence generation and bounding box regression. Similarly, MTL-TabNet visapp23namly also follows the encoder-decoder architecture but contains three decoders for the cell box regression, cell content recognition, and HTML sequence generation. DRCC shen2023divide argues that the error accumulation problem degrades the performance of image-to-sequence TSR models, especially when the input image is large. Therefore, DRCC proposes a two-step decoder architecture, which first decodes the input image into rows and then decodes the rows in cell sequences. VAST huang2023improving pays more attention to the imprecise bounding boxes of table cells and proposes a Coordinate Sequence Decoder to improve the model’s ability to generate accurate bounding boxes and introduces a visual-alignment loss to align the visual and structural information. To sum up, this type of method is usually based on the encoder-decoder architecture and can be trained end-to-end without using post-processing methods.
On the other hand, detection-based models usually define the problem as detecting different table components and applying a post-processing method to reconstruct table structures. DeepTabStR siddiqui2019deeptabstr proposes to detect columns and rows to obtain the table cells. However, DeepTabStR ignores the row/column span in the tables, which means that it cannot recover the hierarchical structures of tables. TableStrRec Fernandes2023TableStrRecFF extends the DeepTabStR, defining four types of table components: regular columns, irregular columns, regular rows, and irregular rows. Then, the spanning cells across multiple columns can be inferred from the difference between the regular and irregular columns when they are overlapped, and the spanning cells across multiple rows can be inferred from regular and irregular rows similarly. PubTables-1M smock2022pubtables is another typical detection-based approach that defines six table components: table, column, row, spanning cell, projected row header, and columns header, in which projected row header and columns header are for the function analysis and other components can be used to reconstruct the complex table structure.
At last, graph-based methods usually apply either detection or segmentation methods to obtain the locations of table cells and further build the relation among table cells. For instance, TGRNet xue2021tgrnet formulates the cell location detection and cell logical location prediction jointly in a multi-task architecture, which is modularized by a segmentation based method and graph convolutional network (GCN), respectively. Similarly, TSRNet li2022table proposes a unified GNN-based approach modeling table detection and table structure recognition tasks together. More specifically, TSRNet also employs a semantic segmentation module to extract primitive regions, then applies k-nearest neighbors and line-of-sight neighbors to construct the graph and further classify the graph nodes and edges to filter the noise regions, merge, and build relations. In contrast, LGPMA qiao2021lgpma proposes a Local Pyramid Mask Alignment Module and Global Pyramid Mask Alignment Module to localize table cells, which are formulated as detection and segmentation problems and can be implemented by MaskR-CNN he2017mask . To construct the structure of the table, LGPMA further proposes a pipeline of cell matching, empty cell searching, and empty cell merging using the Maximum Clique Search algorithm and rule-based methods. Besides building graph explicitly, some studies tensmeyer2019deep ; zhang2022split ; nguyen2023formerge predict the table grids or separators first, and then merge grid elements, which are also treating grid elements as graph nodes. SPLERGE tensmeyer2019deep is a typical method following this strategy consisting of a Split Model and Merge Model, in which the Split Model consists of a Row Projection Network and a Columns Projection Network to obtain the table grid, and the Merge Model is used to merge the grid cells. Similarly, SEM zhang2022split employs a segmentation model to segment columns and rows and generate the table grid with a post-processing method. After the table grid is obtained, SEM introduces an Embedder network to extract and fuse the features from textual and visual modalities. A Merger network takes the fused features from Embedder as inputs to merge the grid elements. TSRFormer-DQ-DETR WANG2023109817 leverages a DETR  carion2020end based separation line prediction model, termed DQ-DETR, to predict the reference points on separation lines, followed by a Relation Network based cell Merging module to merge grid elements.
3 Proposed Method
In this section, we first introduce the overall architecture of the proposed method, then give the details of the proposed methods, including tuning aspect ratios for the anchor boxes generation, refactoring the multi-label definition, and the proposed Spatial Attention Module.
3.1 Overall Architecture
Our proposed solution is based on Cascade R-CNN Cai2018CascadeRD , and the overall architecture is shown in Figure 4. It mainly consists of a backbone network, a region proposal network (RPN), and a series of cascade heads. The backbone network is used to extract features of the input images, which is often followed by Feature Pyramid Network (FPN) lin2017feature to extract and fuse features from different scales. The extracted features, together with the region proposals generated by the RPN, are fed into the first cascade head for the classification and regression tasks, and the regression results would be the inputs of the subsequent cascade head, as shown in Figure 4.
Since there are multiple cascade heads, all the outputs of these cascade heads are used to calculate the loss. Moreover, the final loss of the model can be defined as the sum of these Cascade Heads loss and the RPN loss, as defined by Equation 1 where $N$ is the number of cascade head. It is worth mentioning that we follow the most popular Cascade R-CNN model to use three cascade heads in our model.
$\mathcal{L}=\mathcal{L}_{rpn}+\sum_{i=1}^{N}(\mathcal{L}^{i}_{cls}+\mathcal{L}^{i}_{reg})$ (1)
$e_{cls}=PROJ(ROI\_POOL(BACKBONE(x),b))$ (2)
$e_{reg}=PROJ(ROI\_POOL(BACKBONE(x),b))$ (3)
Figure 4: Overall architecture of Cascade R-CNN.
As shown in Figure 4, there are a REG Head and a CLS Head in each cascade head for the regression and classification tasks, respectively. The input features of these REG Heads and CLS Heads $e_{cls},e_{reg}$ are extracted by applying ROI Pooling operations to the features from Backbone Network with the proposal boxes $b$, which can be defined by Equations 2 and  3. $PROJ,ROI\_POOL$, and $BACKBONE$ are the projection layer, ROI Pooling operation, and the backbone network. Therefore, for a trained model, the features of the CLS Heads $e_{cls}$ are determined by the input image $x$ and the proposal boxes $b$, meaning that a single proposal box cannot be classified into multiple classes because CLS Heads are not multi-label classifiers. As aforementioned, the TSR definition of this study is a multi-label detection problem, we refactor the problem formulation by generating pseudo-classes which will be discussed in Section 3.3. Besides, we also improve the proposal generation of the RPN network by tuning the aspect ratios, and feature generation of the backbone network by incorporating deformable convolution and Spatial Attention, which will be discussed in sections 3.2 and  3.4, respectively.
3.2 Aspect Ratio Tuning
Aspect Ratios are the parameters used in the RPN to generate anchor boxes for the two-stage detectors, which control the shape of the generated anchor boxes. Popular implementations of Cascade R-CNN, such as Detectron2 wu2019detectron2 , usually use 0.5, 1.0, and 2.0 as default values, which can work well for detecting common objects, such as the objects in COCO lin2014microsoft dataset. However, in the context of TSR, the range of aspect ratios is much larger because of the shape of the table components. For example, as shown in Figure 1(c), the aspect ratio of the last row at the table bottom is $20:1084=0.01845$, much smaller than the popular value $0.5$. Therefore, adjusting the aspect ratios to the TSR problem is important. Without proposing fancy new modules to select suitable values, we simply select the values based on the statistics of the training sets. Taking the FinTabNet dataset as an example, the aspect ratios of the defined components are shown in Figure 5. The maximum value is 140, far larger than the popular choices in common object detection. Besides, the majority of aspect ratios in Figure 5 are in the range between 1 and 60. Therefore, we extend this parameter for our proposed TSRDet as [0.0125, 0.025, 0.0625, 0.125, 0.25, 0.5, 1.0, 2.0, 4.0, 8.0, 16, 40, 80]. A further detailed parameter table is provided in Section 4. It is worth mentioning that when an aspect ratio is less than 1, its multiplicative inverse is applied to count the number of aspect ratios in Figure 5. We also didn’t further fine-tune this parameter through validation, which means that they might not be optimal. But this parameter has improved the model performance by around 2.7% as shown in Section 4.3.
Figure 5: Statistics of aspect ratio values of FinTabNet training set. When a aspect ratio is less than 1, then its multiplicative inverse is applied to count the number of aspect ratios.
3.3 Pseudo-class Generation
As mentioned earlier, this study also formulates the TSR problem as detecting six components, namely Table, Column, Row, Spanning Cell, Project Row Header and Column Header, as shown in Figure 1(c) and Figure 3. In the study in smock2022pubtables , all these six components are treated independently. Formally, the ground truth $\mathbf{Y}$ of this definition for each image is a set of tuples containing bounding boxes and their corresponding labels, as defined by Equation 5 where $b_{i},c_{i}$ are the $i$th bounding box and class, and values from 00 to $5$ are the Table, Column, Row, Spanning Cell, Project Row Header and Column Header, respectively. This definition is a multi-label problem for the classification task of a detection model, because some table components, such as Column Header and Row can share identical bounding boxes, as shown in Figure 3. This multi-label definition is challenging for typical two-stage detectors, because the classifiers in these detectors are designed for single-label classification.
$Y=\{(b_{i},c_{i})\}_{i=1}^{N},c_{i}\in\{0,1,2,3,4,5\},\forall i\neq j,(c_{i}\neq c_{j})\land((b_{i}=b_{j})\lor(b_{i}\neq b_{j}))$ (4)
Therefore, in this study, considering the observation that the defined Project Row Headers are all Rows at the same time, we only keep the Project Row Headers samples during the training. Since some Column Headers can share identical bounding boxes with corresponding Rows, we derive a pseudo class for these overlapped samples and remove the original overlapped samples. Therefore, during the training stage, we refactor the ground truth for each image to the regular single-label classification, as defined by Equation 4 where values 00 to $6$ are the Table, Column, Row, Spanning Cell, Project Row Header, Column Header and the Pseudo Class, respectively. During the testing stage, the results of Project Column Header are duplicated once to generate their corresponding prediction Rows, and the results of the pseudo-class are duplicated twice to generate the corresponding prediction Rows and Headers, so that we can still follow the formulation defined by Equation 4 to evaluate the model performance.
$Y=\{(b_{i},c_{i})\}_{i=1}^{N},c_{i}\in\{0,1,2,3,4,5,6\},\forall i\neq j,(c_{i}\neq c_{j})\land(b_{i}\neq b_{j})$ (5)
3.4 Spatial Attention and Deformable Convolution
As discussed in Section 1, building long dependencies for detecting the defined components is important. Inspired by the recent studies using large convolution kernels ding2022scaling ; liu2022convnet , a Spatial Attention Module has been proposed, as shown in Figure 6. For the design of the Spatial Attention Module, we use a similar architecture with MSCA guo2022segnext containing multiple branches and large kernel convolutions and use spatial and depthwise separable convolution chollet2017xception ; howard2017mobilenets to reduce the number of parameters. More specifically, for the spatial separable convolution, we use a pair of $7*1$ and $1*7$ kernels to replace a typical $7*7$, use the pair of $11*1$ and $1*1$ 11, and the pair of $21\times 1$ and $1\times 21$ to replace $11*1$ 11 and $21*21$ kernels, respectively. For the depthwise separable convolution, we applied the convolution on each channel of the feature maps independently. Then, the outputs of the three branches are concatenated together as the input of a convolution layer with $1\times 1$ kernel to make the channel dimension the same as the inputs. The proposed Spatial Attention Module can be easily inserted into the backbone network between two blocks because they do not change the feature shapes. For example, for a typical backbone network implemented by ResNet he2016deep containing a STEM block and four Residual Blocks, as shown in Figure 6, the Spatial Attention module can be inserted after the last three Residual Blocks to generate the spatial attention, then the spatial attention can be applied to the original outputs of each Residual Block by Element-wise Multiplication. It is worth mentioning that the Spatial Attention Modules shown in Figure 6 have independent trainable parameters, and all the feature maps are padded correspondingly to keep the size of the feature maps.
Figure 6: Architecture of proposed Spatial Attention Module. A ResNet backbone consists of a STEM Block and four stages of Residual Block. Our proposed Spatial Attention Module are inserted between the blocks of the backbone to build long dependencies.
On the other hand, deformable convolution dai2017deformable has been demonstrated to help improve the detection performance on the document image detection tasks by many studies fernandes2023tablestrrec ; siddiqui2018decnt ; mondal2023dataset . Deformable convolution uses a learnable grid offset to sample the grid points from the feature map and apply the convolution operation, as defined by Equation 6,
$\mathbf{z}_{p_{0}}=\sum_{p_{n}\in R}{w(p_{n})x(p_{0}+p_{n}+\Delta p_{n})}$ (6)
where $p_{0}$ is the location on the output feature map $\mathbf{z}$, $p_{n}$ is the $n$th grid point in grid $R$, and $\Delta p_{n}$ is the $n$th learnable offset. Since the offset $\Delta p_{n}$ applied to the deformable convolution is usually obtained by a regular convolution with small kernels, such as a 3 * 3 kernel, it can only improve the local feature instead of building long dependencies. Therefore, deformable convolution is useful for improving the detection performance but can contribute little to exploring the structures of the table. In this study, we apply the proposed Spatial Attention module and deformable convolution to build the long-dependent and local features together.
4 Experiments and Analysis
4.1 Datasets and Experimental Settings
We utilize four datasets in this work, including SciTSR chi2019complicated , FinTabNet zheng2020global , PubTabNet zhong2020image and PubTables1M smock2022pubtables . As discussed in the study in smock2023aligning , FinTabNet and SciTSR datasets contain noise annotations that are harmful for the model performance. Therefore, we use their cleaned versions proposed in the study in smock2023aligning . All the four datasets consist of only the table regions of each image. SicTSR dataset is collected from academic publications, which contains 7453, 1034 and 2134 samples for the training, validation and testing. PubTables1M dataset is a large scale dataset for the TSR problem collected from the PMCOA corpus, containing 758849 training samples, 94959 validation samples, and 93834 testing samples. Since the PubTabNet dataset does not provide original PDF files, we cannot process it to make detection annotations. Besides, its testing is not public available. Therefore, we use its validation set to evaluate the model trained with the PubTable1M dataset. Following the study in smock2023aligning , we use the code base in smock2021tabletransformer to process the datasets and align the formats of all the three datasets. FinTabNet is also a large dataset widely used for the TSR problem, containing 78537, 9289 and 9650 samples for the training, testing and validation, respectively. FinTabNet is collected from the annual reports of companies, making its data source different from other two datasets. Table 1 summarizes the datasets used in this study for the model evaluation.
Table 1: Summary of datasets.
Dataset Train Validation Test
SciTSR chi2019complicated 7,453 1,034 2,134
FinTabNet zheng2020global 78,537 9,650 9,289
PubTabNet zhong2020image 500,777 9,115 -
PubTables1M smock2022pubtables 758,849 94,959 93,834
Since the TSR problem in this study is formulated as an object detection problem, we use both detection metrics and cell-level TSR metric for the model evaluation. For the detection metrics, we employ the widely accepted COCO metrics lin2014microsoft , namely mean Average Precision (mAP), $AP50$, $AP75$, $AP_{s}$, $AP_{m}$, $AP_{l}$, and object specific AP scores. More specifically, $AP50$, $AP75$ are the APs using 0.50 and 0.75 as IoU thresholds, respectively. $AP_{s}$, $AP_{m}$ and $AP_{l}$ are the APs of different size of target objects, which is defined by Equation 7.
[…]
For the TSR metric, we choose structure-only Tree-Edit-Distance-Based Similarity(TEDS) zhong2020image , which is firstly introduced in the study to overcome the drawbacks of adjacency relation metrics hurst2003constraint , and can be defined as Equation 8.
$TEDS(T_{a},T_{b})=1-\frac{EditDist(T_{a},T_{b})}{max(\mid T_{a}\mid,\mid T_{b}\mid)}$ (8)
where EditDist is the tree-edit distance, and $T$ is the number of nodes in the tree. It is worth mentioning that we use structure-only TEDS in the study, which only consider the HTML tags without extracting their contents to avoid the influence of OCR tools. It is worth mentioning that the testing samples can be categorized into Sim. and Com. groups based on whether they have cells spanning across multiple columns and rows.
4.2 Implementation Details and Experimental Results
To verify the effectiveness of our proposed solution, we include all three types of state-of-the-art methods as discussed in Section 2. For the detection-based methods, Cascade R-CNN Cai2018CascadeRD , Deformable-DETR zhu2020deformable and Sparse R-CNN sun2021sparse as benchmark models, in which Cascade R-CNN Cai2018CascadeRD is also the based model of the proposed methods, Deformable-DETR and Sparse R-CNN are two state-of-the-art transformer based models. For the image-to-sequence models, EDD zhong2020image , GTE zheng2020global , TableFormer nassar2022tableformer , TableMaster ye2021pingan , VAST huang2023improving and MTL-TabNet visapp23namly are included. TSRFormer-DQ-DETR WANG2023109817 is a state-of-the-art model following the pipeline of detecting separation lines and then merging cell grids, as discussed in Section 2.
We implement Cascade R-CNN and our proposed TSRDet based on the Detection2 wu2019detectron2 , build the Deformable-DETR based on detrex ren2023detrex , train the Sparse R-CNN with their official codebase. For the Deformable-DETR and Sparse R-CNN, we use the default parameters. For the Cascade R-CNN baseline, we align the number of regional proposals and the batch normalization method to the TSRDet, as shown in Table 2. All these detection models are using ResNet50 he2016deep pre-trained with ImageNet deng2009imagenet as the backbone network. We also re-train TableMaster ye2021pingan with the FinTabNet dataset based on its official code base. For the implementation of the proposed TSRDet, aspect ratios in the anchor box generation are set as [0.0125, 0.025, 0.0625, 0.125, 0.25, 0.5, 1.0, 2.0, 4.0, 8.0, 16, 40, 80], and other key parameters are summarized in Table 2. Notably, to calculate the structure-only TEDS, we use the scripts provided by study in smock2021tabletransformer to generate the HTML sequences from the detected components. All the models are trained with 240, 120, and 60 epochs for the SciTSR, FinTabNet, and PubTables1M datasets, respectively.
The experimental results regarding the COCO metrics and structure-only TEDS are shown in Tables 7,  34,  5 and  6, which can demonstrate the superiority of the proposed solution. For the SciTSR dataset, the proposed TSRDet can improve the baseline Cascade R-CNN by 19.32% regarding the structure-only TEDS, outperforming Deformable-DETR and Sparse R-CNN. When it comes to COCO metrics, the mAP of the proposed TSR is as good as Deformable-DETR, outperforming other benchmark models. Similarly, the proposed TSR can outperform benchmark models regarding both COCO metrics and structure-only TEDS on the FinTabNet and PubTables1M datasets. It is worth mentioning that FinTabNet does not provide original PDF files, making it hard to generate detection annotations. Therefore, the model performance reported in Table 6 is calculated using the model trained with the PubTable1M dataset. Even though the PubTable1M and PubTabNet datasets have misalignments regarding the ground truth HTML sequences, the proposed method still shows competitive performance compared with other state-of-the-art methods, as shown in Table 6. Figure 7 shows a prediction sample and its generated HTML sequence after post-processing, demonstrating the capacities of the proposed solution.
Besides demonstrating the effectiveness of the proposed TSRDet, there are some important observations from the experimental results. First, structure-only TEDS is not well aligned with COCO metrics. For example, even though the baseline Cascade R-CNN model can only reach 87.49% regarding the structure-only TEDS on the FinTabNet dataset, its detection performance is only slightly worse than other models. Deformmable-DETR has better mAP, but worse structure-only TEDS than Sparse R-CNN model on the FinTabNet dataset. Second, benchmark models, especially Sparse R-CNN also have very promising performance even though it is not as good as our proposed solution. The discussions and analysis to these observations will be in Section 4.4.
Table 2: Key training parameters of the proposed model. MAX_ITER and STEPS are for the FinTabNet dataset as examples.
Parameter Value Description
RESNETS.NORM nnSyncBN Batch Normalization for the Backbone Network
MAX_ITER 112,500 total number of mini-batch
STEPS 84,375 the mini-batch to apply the learning rate schedule
SCHEDULER MultiStepLR the scheduler to change the learning rate
NMS_THRESH 0.9 non-maximum suppression threshold
PRE_NMS_TOPK_TRAIN 4000 RPN proposals to keep before applying NMS
PRE_NMS_TOPK_TEST 2000 RPN proposals to keep before applying NMS
POST_NMS_TOPK_TRAIN 4000 RPN proposals to keep after applying NMS
POST_NMS_TOPK_TEST 2000 RPN proposals to keep after applying NMS
DEFORM_ON_PER_STAGE [True, True, True, True] whether to use deformable convolution in each backbone stage
Table 3: Experimental results on SciTSR dataset with structure-only TEDS score. Sim. means the tables without spanning cells and Com. represents the tables with spanning cells.
Model TEDS-struc.(%)
Sim. Com. All
Cascade R-CNN 77.31 84.74 79.09
Deformable-DETR 98.17 94.59 97.30
SparseR-CNN $\mathbf{99.08}$ 95.92 98.30
TSRDet(Ours) 98.59 $\mathbf{97.88}$ $\mathbf{98.41}$
Table 4: Experimental results on FinTabNet dataset with structure-only TEDS score. Sim. means the tables without spanning cells and Com. represents the tables with spanning cells.
Model TEDS-struc.(%)
Sim. Com. All
EDD zhong2020image 88.40 92.08 90.60
GTE zheng2020global - - 87.14
GTE${}^{(FT)}$ zheng2020global - - 91.02
TableFormer nassar2022tableformer 97.50 96.00 96.80
TableMaster ye2021pingan 98.36 98.28 98.32
VAST huang2023improving - - 98.63
MTL-TabNet visapp23namly 99.07 98.46 98.79
TSRFormer-DQ-DETR WANG2023109817 - - 98.40
Cascade R-CNN 82.17 92.50 87.49
Deformable-DETR 98.08 97.54 97.81
Sparse-RCNN 98.36 97.91 98.13
TSRDet(Ours) $\mathbf{99.08}$ $\mathbf{99.02}$ $\mathbf{99.05}$
Table 5: Experimental results on PubTables1M dataset with structure-only TEDS score. Sim. means the tables without spanning cells and Com. represents the tables with spanning cells.
Model TEDS-struc.(%)
Sim. Com. All
Cascade R-CNN 82.73 85.21 83.78
Deformable-DETR 97.54 93.14 95.73
SparseR-CNN 99.04 95.90 97.72
TSRDet(Ours) $\mathbf{99.19}$ $\mathbf{97.66}$ $\mathbf{98.55}$
Table 6: Experimental results on PubTabNet validation set with structure-only TEDS score. Sim. means the tables without spanning cells and Com. represents the tables with spanning cells. The proposed model is trained with PubTable1M dataset, while the benchmark models are trained with PubTabNet dataset.
Model TEDS-struc.(%)
Sim. Com. All
EDD zhong2020image 91.10 88.70 89.90
GTE zheng2021global - - 93.01
LGPMA qiao2021lgpma - - 96.70
RobustTabNet ma2023robust - - 97.00
TSRNet li2022table - - 95.64
VAST huang2023improving - - 97.23
TableFormer nassar2022tableformer 98.50 95.00 96.75
MTL-TabNet visapp23namly 99.05 96.66 97.88
TSRDet(Ours) 96.99 94.99 96.58
Table 7: Experimental results with Average Precision.
Dataset Model $AP$ $AP50$ $AP75$ $AP_{s}$ $AP_{m}$ $AP_{l}$ table column row spanning cell projected row header column header SciTSR Cascade R-CNN 93.89 95.27 94.80 95.81 93.89 92.96 98.96 98.63 96.33 88.58 83.80 97.01 Deformable-DETR 96.28 97.39 97.01 96.75 96.55 96.07 98.96 98.63 97.26 93.84 90.86 98.15 Sparse-RCNN 94.78 96.17 95.48 95.49 95.07 90.08 98.98 98.30 97.93 88.06 86.92 98.49 TSRDet(Ours) 96.28 96.79 96.57 99.01 96.42 95.65 98.97 99.25 98.57 95.30 87.06 98.50 FinTabNet Cascade R-CNN 95.23 97.53 96.90 87.32 95.31 93.08 99.00 96.69 96.96 84.43 96.63 97.64 Deformable-DETR 96.68 98.42 97.98 75.17 95.53 95.58 99.00 97.55 96.95 91.91 96.62 98.04 Sparse-RCNN 96.38 98.37 97.69 62.11 96.22 95.86 99.01 97.79 97.84 88.39 97.29 97.97 TSRDet(Ours) 97.50 98.33 98.09 91.60 97.40 97.15 99.01 98.83 97.99 94.62 96.61 97.93 PubTables1M Cascade R-CNN 93.40 95.38 94.76 85.75 93.32 92.57 99.01 98.76 87.56 82.18 95.81 97.11 Deformable-DETR 94.82 97.43 96.79 78.33 92.55 94.48 98.99 97.89 95.84 85.04 95.43 95.74 Sparse-RCNN 96.46 98.14 97.60 84.25 95.73 96.45 99.00 98.42 98.03 87.85 97.91 97.57 TSRDet(Ours) 97.72 98.26 98.04 94.76 97.43 97.33 99.01 98.99 98.41 94.21 97.88 97.85
(a) A prediction sample after post-processing.
(b) The generated HTML sequence after post-processing.
Figure 7: A sample of prediction result from the FinTabNet testing set.
4.3 Ablation Study
In this section, we conduct experiments on the FinTabNet dataset to demonstrate the effectiveness of our applied methods, including applying aspect ratio tuning, pseudo-class generation, deformable convolution and spatial attention. The experimental results are shown in Tables 8 and 9, in which Asp_Ratio Tuning, Pseudo-Class Gen, DEFORM, and S_Attn are shorts for applying aspect ratio tuning, pseudo-class generation, deformable convolution, and spatial attention, respectively. We use the default implementation of Cascade R-CNN in Detectron2 wu2019detectron2 as the baseline model. Even though this default implementation can reach 95.06% regarding the mAP, its overall structure-only TEDS only reaches 82.70%. After tuning the aspect ratios for the anchor generation, the structure-only TEDS is increased to 90.23%, even though the mAP is only increased from 95.06% to 95.54%. Other ablation results show similar trends compared with the baseline model. It is worth mentioning that simply applying deformable convolution without other methods can increase the detection performance significantly, but lead to worse structure-only TEDS score, , which is a similar phenomenon aforementioned in Section 4.2. This is because deformable convolution can make model fit better to the extra spaces that are not necessarily useful for recovering the table structure.
Table 8: Ablation study results on FinTabNet dataset with structure-only TEDS score. Asp_Ratio Tuning, Pseudo-Class Gen, DEFORM, and S_Attn are shorts for applying aspect ratio tuning, pseudo-class generation, deformable convolution, and spatial attention.
Model Asp_Ratio Tuning Pseudo-Class Gen DEFORM S_Attn TEDS-struc.(%)
Sim. Com. All
Cascade R-CNN 82.17 92.50 87.49
Ablation 1 81.45 87.11 84.35
Ablation 2 84.27 95.80 90.23
Ablation 3 95.17 98.63 96.95
Ablation 4 96.44 99.14 97.83
Ablation 5 96.95 98.75 97.88
TSRDet(Ours) 99.08 99.02 99.05
Table 9: Ablation study results regarding Mean Average Precision. The model names are aligned with models in Table 8.
Model $AP$ $AP50$ $AP75$ $AP_{s}$ $AP_{m}$ $AP_{l}$ table column row spanning cell projected row header column header Cascade R-CNN 95.23 97.53 96.90 87.32 95.31 93.08 99.00 96.69 96.96 84.43 96.63 97.64 Ablation 1 97.22 98.03 97.90 90.11 96.72 96.76 99.00 98.95 96.16 94.98 96.01 98.19 Ablation 2 95.54 97.54 96.91 87.43 95.79 94.04 99.00 97.04 97.64 84.84 96.67 98.02 Ablation 3 95.51 97.56 96.94 88.43 95.52 93.76 99.00 97.31 97.87 84.74 96.82 97.28 Ablation 4 97.83 98.37 98.13 91.91 97.65 97.58 99.00 98.96 98.33 95.78 96.98 97.93 Ablation 5 96.97 97.84 97.58 90.32 96.88 96.21 99.00 98.83 98.03 91.97 96.58 97.37 TSRDet(Ours) 97.50 98.33 98.09 91.60 97.40 97.15 99.01 98.83 97.99 94.62 96.61 97.93
4.4 Observations and Analysis
As mentioned earlier in Sections 4.2 and  4.3, there are some important observations from the experimental results that can demonstrate the key design aspects included in this study. In this section, we further discuss and explain these important observations.
4.4.1 Metrics misalignment
The experimental results in Sections 4.2 and  4.3 show that a model with better mAP performance does not necessarily also have better structural TEDS, which means that the structure-only TEDS and the COCO metrics are not aligned. For example, in Table 7, both the Deformable-DETR and our proposed TSRDet can reach 96.28% regarding mAP on the SciTSR dataset, which is better than that of Sparse R-CNN. However, when it comes to structure-only TEDS, as shown in Table 3, both TSRDet and Sparse R-CNN can perform better than Deformable-DETR. COCO metrics evaluate the similarity between the prediction and ground truth boxes with IoU thresholds. In contrast, the structure-only TEDS evaluates the structure difference between the detected components and the ground truth components because the post-processing method uses structure information, such as the overlaps between columns and rows, to generate table cells. In the context of detection-based TSR, the defined components often contain large white areas that can influence the COCO metrics, but cannot influence the structure. Figure 8 shows the columns’ ground truth of a sample from the FinTabNet dataset. As shown in Figure 7(a), the minimum boxes required to infer the correct architecture are much smaller than the ground truth boxes. Figure 7(b) and  7(c) show two prediction samples after post-processing using Ablation 1 and Ablation3 models. Ablation 1 model with deformable convolution can better fit the extra white areas to improve the detection performance, but it is not necessarily helpful for the structure.
(a) Comparison of the ground truth bounding box and the minimum bounding box for the structure. We only include the columns’ boxes for simplicity.
(b) A sample prediction result of the Ablation 1 model after post-processing. The mAP and structure-only TEDS are 97.22 and 84.35, respectively. We only include the columns’ predictions for simplicity.
(c) A sample prediction result of the Ablation 3 model after post-processing. The mAP and structure-only TEDS are 95.51 and 96.95, respectively. We only include the columns’ predictions for simplicity.
Figure 8: Comparison of results from Ablation1 and Ablation3 models. Even though Ablation 1 can achieve better detection performance, its performance regarding structure-only TEDS is much lower than that of Ablation 3 model.
4.4.2 Comparison with Sparse R-CNN
Besides the metrics misalignment, our other important observation is that the Deformable-DETR and Sparse R-CNN also achieved promising results even though they are not as good as the proposed method. Therefore, using Sparse R-CNN as an example, we further analyze the difference between our proposed TSRDet and Sparse R-CNN and discuss how Sparse R-CNN implicitly deals with the multi-label definition.
Sparse R-CNN mainly consists of a backbone network, learnable region proposal, and a series of Dynamic Heads containing Dynamic Instance Interactive Head. Therefore, for a input image $x$ with initialized region proposal feature embedding $e_{pro}$ and corresponding proposal boxes $b$, the inputs to the $head_{cls}$ and $head_{reg}$ in the first Dynamic Heads can be generated by Equations 910,  11 and 12, in which $BACKBONE,ROI\_POOL,SELF\_ATTN$ and $PROJ$ are the backbone network, ROI_Pooling, self attention and projection layers. Since Sparse R-CNN also utilizes a cascade architecture, the $N$th Dynamic Head takes the outputs of $(N-1)$th Head’s regression result and $e_{dyn}$ as the proposal boxes and proposal feature embedding. Since Sparse R-CNN has multiple ways to initialize the region proposal boxes, such as Random boxes, Image Size boxes, and Grid boxes, we follow the setting of Section 4 to assume the region proposal boxes are initialized with Image Size boxes, which means that for an input image $x$, its proposal boxes $b$ can be treated as constants. Therefore, the classification and regression results are determined by the learned region proposal feature embedding $e_{pro}$, which can differ even when their corresponding boxes are identical. Therefore, Sparse R-CNN can naturally deal with the multi-label definition for the TSR task.
$e_{roi}=ROI\_POOL(BACKBONE(x),b)$ (9)
$e_{dyn}=DYN\_CONV(SELF\_ATTN(e_{pro},e_{pro},,e_{pro}),e_{roi})$ (10)
$e_{cls}=PROJ_{cls}(e_{dyn})$ (11)
$e_{reg}=PROJ_{reg}(e_{dyn})$ (12)
By contrast, for the classification head of Cascade R-CNN, their inputs are features extracted by applying RoI Pooling operation, as defined by Equation 2. In other words, the input features of the classification head are determined by the bounding boxes once the model is trained. Therefore, there can only be one classification result for a proposal box, making Cascade R-CNN challenging to deal with the multi-label definition for the TSR task.
5 Conclusion and Future Work
In this study, we propose TSRDet, a tailored Cascade R-CNN based solution for the TSR task, directly detecting six types of table components. The experimental results show that the proposed solution outperforms different types of state-of-the-art models, including image-to-sequence and graph-based models. Considering the definition of six components, a bounding box can simultaneously be a Row, Projected Row Header, and Column Header, which means it is a multi-label classification. We refactor the multi-label classification to regular single-label classification by generating pseudo-classes. Besides, the proposed Spatial Attention module can be inserted into the backbone network to build long-range dependencies, so deformable convolution is applied to further improve the model performance. Since the proposed method is detection-based and focuses on well-formatted, visually rich documents, one major limitation is that it may fail to deal with irregular tables, such as rotated and distorted tables. Integrating instance segmentation with detection models is included in our future research agenda to deal with irregular tables because instance segmentation can handle the irregular shape and be guided by the bounding boxes.
Acknowledgements
This work was supported in part by Mathematics of Information Technology and Complex Systems (Mitacs) Accelerate Program and Lytica Inc.